OpenStudy (republican31):

Inka can mow a lawn in 2.5 hours, and Phil can mow the same lawn in 3.5 hours. How many hours will it take them, working together, to mow the lawn? Round your answer to the nearest tenth of an hour. Is the proper way to set it up x/2.5 + x/3.5 = 1?

OpenStudy (sepeario):

Inka does 1/2.5 of the job an hour, and Phil does 1/3.5, so together in h hours ∙ ∙ h( 1/2.5 + 1/3.5) = 1 whole job ∙ ∙ h/2.5 + h/3.5 = 1 ∙ ∙ ∙ ∙ ∙ ∙ ∙ ∙ ∙ ∙ ∙ multiply by 2.5 and 3.5 ∙ ∙ 3.5h + 2.5h = 2.5 • 3.5 ∙ ∙ 6h = 8.75 ∙ ∙ h = 1.4583333 hr ≈ 1.5 hr

OpenStudy (anonymous):

no, it's 1/2.5 + 1/3.5 = 1/x, where x is the time it takes both of them to finish

OpenStudy (republican31):

I'm still confused. Can we do this a little more step by step?